OpenStudy (anonymous):

I NEED HELP WITH 10 QUESTIONS IN MATH I WILL FAN YOU AND GIVE YOU A MEDAL PLEASE HELP!!!

OpenStudy (anonymous):

what r the questions

OpenStudy (anonymous):

#1 what is the quotient: (2x^2+10x+12) divided by (x+3)? a= 2x+6 b-2x-6 c=2x+4 d=2x-4,r=1

OpenStudy (anonymous):

#2 what is the quotient: (6x^4-15x^3-2x^2-10x-4) divided by (3x^2+2) a= 2x^2-5x+2 b= 2x^2+5x-2 c=2x^2-5x-2 D= 2x^2+5x+2

OpenStudy (anonymous):

https://www.mathway.com/ my life saver!!!!!

OpenStudy (anonymous):

its amazing! I use it all the time

OpenStudy (anonymous):

okay thanks :D